Project

General

Profile

Actions

2018-05-29-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

15th May 2018; 15:30 BST
RAL, R1 1st Floor North, CR22

Present: KL, SB, DR, CW

Notes

  1. Introduction: KL
    • Routine meeting. KL thanked all present for accommodating the change of time caused by his teaching commitment.
  2. Notes of previous meeting and status of actions: All
    • Notes of the previous meeting will be posted and comments will be noted at the next meeting. Status of actions:
      • CW: With others make inventry of what may not be scrapped by ISIS post June 2018 when ISIS take over ownershp of the MICE Hall.
        • Done. CW has circulated an updated draft inventory. People should take note and comment as required.
      • PS: Generate two-pager to close-out common fund for next PSAG.
        • Stands.
  3. Project Manager's report: CW
    • Most issues covered in the Hall inventory. JG is off sick this week and we are not sure when he will be back.
  4. Software and computing status: DR
    • DR and V.Pec are working on TOF bad channel maps. TOF review meetings will resume next week.
    • Geometry:
      • Tracker glue issue is fixed.
      • Survey of the diffuser has been completed. Clarification of the iris thicknesses still needs to be obtained. CR is working with A.Millington.
    • The cuts framework is being tested for inclusion in the trunk.
    • Once a new release has been made, all data will be reprocessed. V.Blackmore has requested a new MC and processing for the emittance-measurement paper.
    • MLCR and computing decommissioning/plan for continued support for analysis:
      • Final steps in removing our dependency on computing in the MLCR are being carried out. Only the PCs for batch processing will remain in the MLCR. A machine to do the a CDB update will also remain live in the MLCR.
    • Future proofing the MICE WWW site:
      • Remains an action on DR. Issue with licence on IIT WWW server has been solved. Agree to migrate mail issues to jiscmail. DR will drive it through.
    • Taking over from HN:
      • Plan has been made and sent to KL and P.Soler to be taken forward. Require to pass information from H.Nebrensky to new person befre the end of June 2018.
  5. Paper status and plans: CR
    • In CR's absense, KL reported that the emittance-measurement paper was now with the referees who are required to comment by Friday 01Jun18. The LH2 paper is also with the authors for final comments. Chris Hunt has agreed to lead the preparation of the long emittance-evolution paper. A plan is expected from him in about a weeks time.
    • Tanaz -- plans for coming to UK? Need to prompt Pavel and Dan.
  6. AOB

Summary of actions:

  • PS: Generate two-pager to close-out common fund for next PSAG.
  • DR: Continue to push towards a persistence plan for the MICE IIT WWW site.
  • DR: Coordinate migration of mail server from IIT to jiscmail.

Back to MEMO:

Back

Back to MIPO:

Back

Documents

Updated by Long, Kenneth almost 5 years ago · 1 revisions